shanorea:

Paul is a salesperson who earns both a salary and a commission on units sold. In one week, he sold 25 units and earned $1,200. In another week, he sold 20 units and earned $1,075. Which of the following slope-intercept form equations models Paul’s weekly pay in terms of units sold?

yomomalikebracobama:

hmmm let me check this out

yomomalikebracobama:

he sold 25 units earned 1,200 the FIRST WEEK another week he sold 20 and earned 1075 so where are the awnser choises?

shanorea:

wait my bad

yomomalikebracobama:

lol

shanorea:

y = 25x + 575 B. y = 25x + 1,200 C. y=-1/25x - 68 D. y=1/25x - 23

shanorea:

nvm